# Hurwitz moduli varieties parameterizing pointed covers of an algebraic curve with a fixed monodromy group

### v3

@@ 1. Introduction
The covers $f:X\to Y$ with restricted monodromy group and the related Galois covers $p:C\to Y$ yield polarized abelian varieties isogenous to abelian subvarieties of the Jacobian variety $J(X)$ [6, 16, 25, 4]. Given a smooth, proper family of such covers one obtains a morphism of its parameter variety to a certain moduli space of polarized abelian varieties, which may be studied by means of the associated variation of the Hodge structures of weight one [13, 3]. This approach to the moduli of polarized abelian varieties was used in [17, 18] where the unirationality of $\mathcal{A}_{3}(1,1,d)$ and $\mathcal{A}_{3}(1,d,d)$ with $2\leq d\leq 4$ was proved by means of families of simply ramified covers of elliptic curves branched in 6 points. The Prym-Tyurin morphism from a Hurwitz space to $\mathcal{A}_{6}$ was used in [1] to prove that every sufficiently general principally polarized abelian variety of dimension 6 is isomorphic to a Prym-Tyurin variety of a cover of $\mathbb{P}^{1}$ of degree 27, branched in 24 points, with monodromy group the Weyl group $W(E_{6})\subset S_{27}$.
@@
We assume the base field is $\mathbb{C}$. Algebraic varieties are reduced, separated, possibly reducible schemes of finite type, points are closed points. Fiber products and pullbacks are those defined in the category of schemes over $\mathbb{C}$. A cover $f:X\to Z$ of algebraic varieties is a finite, surjective morphism. If $Y$ is a smooth, projective, irreducible curve we denote by $Y^{(n)}_{\ast}$ the subset of effective divisors of $Y$ of degree $n$ without multiple points and identify such divisors with their support. If $\Omega\subset Y^{(n)}$, then $\Omega_{\ast}=\Omega\cap Y^{(n)}_{\ast}$. Given an algebraic variety $(X,\mathcal{O}_{X})$ the canonically associated complex space is denoted by $(X^{an},\mathcal{O}_{X^{an}})$ [27]. Its topological space is denoted by $|X^{an}|$. We assume that the homotopy of paths leaves the end points fixed [23, Chapter 2, § 2]. The set of paths homotopic to $\alpha$ is denoted by $[\alpha]$. The product of the paths $\alpha$ and $\beta$ is denoted by $\alpha\cdot\beta$ and it equals the path $\gamma$, where $\gamma(t)=\alpha(2t)$ if $t\in[0,\frac{1}{2}]$, $\gamma(t)=\beta(2t-1)$ if $t\in[\frac{1}{2},1]$. Given a covering space $p:M\to N$ of the topological space $N$, the map $p$ is called topological covering map. Lifting a path $\alpha$ of $N$ from an initial point $z\in M$ the terminal point is denoted by $z\alpha$.
@@ 2. Parameterization of covers with a fixed monodromy group
Throughout the paper $Y$ is a smooth, projective, irreducible curve of genus $g\geq 0$, $n$ is a positive integer, $G$ is a finite group, which can be generated by $12g+n-$ 1 elements, $\Lambda$ is a finite set, $|\Lambda|=d$, on which $G$ acts on the right faithfully and transitively, we choose an element $\lambda_{0}\in\Lambda$ as a marked element, $G(\lambda_{0})=\{g\in G|\lambda_{0}g=\lambda_{0}\}$. We identify $G$ with its image in the symmetric group $S(\Lambda)\cong S_{d}$, where $S(\Lambda)$ acts on the right.
@@
A necessary condition for two covers $f:X\to Y$ and $f_{1}:X_{1}\to Y$ to be equivalent is that they have the same branch locus $D$. Let $y_{0}\in Y\setminus D$. Let $Y^{an}$ be the Riemann surface associated with $Y$. Then $\pi_{1}(Y^{an}\setminus D,y_{0})$ acts on the right on $f^{-1}(y_{0})$ by $x\cdot[\alpha]=x\alpha$ for $\forall x\in f^{-1}(y_{0})$ and $\forall[\alpha]\in\pi_{1}(Y^{an}\setminus D,y_{0})$, and similarly it acts on the right on $f_{1}^{-1}(y_{0})$. Then $f$ is equivalent to $f_{1}$ if and only if there exists a $\pi_{1}(Y^{an}\setminus D,y_{0})$-equivalent bijection $\mu:f^{-1}(y_{0})\to f_{1}^{-1}(y_{0})$. In fact let $X^{\prime}=f^{-1}(Y\setminus D)$, $X_{1}^{\prime}=f_{1}^{-1}(Y\setminus D)$. Then such a bijection yields a covering biholomorphic map $X^{\prime an}\to X^{\prime an}_{1}$ over $Y\setminus D$, which extends to a covering biholomorphic map $X^{an}\to X^{an}_{1}$ over $Y^{an}$. Furthermore this map equals $h^{an}$, where $h:X\to X_{1}$ is a covering isomorphism over $Y$. Abusing notation we will write $\pi_{1}(Y\setminus D,y_{0})$ instead of $\pi_{1}(Y^{an}\setminus D,y_{0})$.
@@
Let Let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$. Let $\overline{U}_{1},\ldots,\overline{U}_{n}$ be embedded closed disks in $Y\setminus y_{0}$ which are disjoint and such that $b_{i}\in U_{i}$ for $\forall i$, where $U_{i}$ is the interior of $\overline{U}_{i}$. For every $i=1,\ldots,n$ let us choose a path $\eta_{i}:I\to Y\setminus\cup_{j=1}^{n}U_{j}$ such that $\eta_{i}(0)=y_{0}$, $\eta_{i}(1)\in\partial\overline{U}_{i}$ and let $\gamma_{i}:I\to Y\setminus D$ be the closed path which starts at $y_{0}$, travels along $\eta_{i}$, then makes a counterclockwise loop along $\partial\overline{U}_{i}$ and returns back to $y_{0}$ along $\eta_{i}^{-}$. Let $f:X\to Y$ be a $d$-sheeted cover of $Y$ unbranched at $y_{0}$. Let $X^{\prime}=X\setminus f^{-1}(D)$, $f^{\prime}=f|_{X^{\prime}}$. The branch locus of $f$ equals $D$ if and only if $f^{\prime}:X^{\prime}\to Y\setminus D$ is unramified and the monodromy $m:\pi_{1}(Y\setminus D,y_{0})\to S(f^{-1}(y_{0}))$ defined by $x\cdot m([\alpha])=x\alpha$ for $\forall x\in f^{-1}(y_{0})$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$ satisfies the condition
@@
Let $m:\pi_{1}(Y\setminus D,y_{0})\to S_{d}$ be a homomorphism and let $G=\operatorname{Im}(m)$. Suppose that $m$ satisfies Condition (1). If one chooses in a different way $\overline{U}_{1},\ldots,\overline{U}_{n}$ and $\eta_{1},\ldots,\eta_{n}$ so as to satisfy the above conditions, then for every $i\in[1,n]$ the new element $[\gamma^{\prime}_{i}]$ belongs to the conjugacy class of $[\gamma_{i}]$ in $\pi_{1}(Y\setminus D,y_{0})$, so $m([\gamma^{\prime}_{i}])$ belongs to the conjugacy class of $m([\gamma_{i}])$ in $G$. Therefore Condition (1) for $m$ is independent of the choice of $\overline{U}_{1},\ldots,\overline{U}_{n}$ and that of $\eta_{1},\ldots,\eta_{n}$.
@@
Let $G\subset S(\Lambda)$ be as in § 2.1. Let $N_{S(\Lambda)}(G)=\{\sigma\in S(\Lambda)|\sigma G\sigma^{-1}=G\}$. Let $f:X\to Y$ and $f_{1}:X_{1}\to Y$ be two covers with the same branch locus $D$. Let $y_{0}\in Y$. Suppose that there are bijections $\varepsilon:\Lambda\to f^{-1}(y_{0})$, $\varepsilon_{1}:\Lambda\to f_{1}^{-1}(y_{0})$ and epimorphisms $m:\pi_{1}(Y\setminus D,y_{0})\to G$, $m_{1}:\pi_{1}(Y\setminus D,y_{0})\to G$ such that $\varepsilon(\lambda m([\alpha])=\varepsilon(\lambda)\alpha$ and $\varepsilon_{1}(\lambda m_{1}([\alpha])=\varepsilon_{1}(\lambda)\alpha$ for $\forall\lambda\in\Lambda$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$. Then $f$ is equivalent to $f_{1}$ if and only if there exists a $\sigma\in N_{S(\Lambda)}(G)$ such that $m_{1}=\sigma m\sigma^{-1}$. The covering isomorphism $h:X\to X_{1}$ is unique if and only if the centralizer of $G$ in $S(\Lambda)$ is trivial, $Z_{S(\Lambda)}(G)=\{1\}$.
@@
The covers $f:X\to Y$ and $f_{1}:X_{1}\to Y$ are equivalent if and only if there exists a $\pi_{1}(Y\setminus D,y_{0})-$equivariant bijection $\mu:f^{-1}(y_{0})\to f_{1}^{-1}(y_{0})$ (cf. § 2.3). Let $\mu$ be such a bijection. Let $s:\Lambda\to\Lambda$ be the bijection, which makes the following diagram commutative
(2) ${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon}$$\scriptstyle{s}$${f^{-1}(y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varepsilon_{1}}$${f_{1}^{-1}(y_{0})}$
@@
$\begin{split}&\mu(x\alpha)=\mu(\varepsilon(\lambda)\alpha)=\mu(\varepsilon(\lambda m([\alpha])))=\varepsilon_{1}(\lambda m([\alpha])\sigma^{-1})\\ &\mu(x)\alpha=\mu(\varepsilon(\lambda))\alpha=\varepsilon_{1}(\lambda\sigma^{-1})\alpha=\varepsilon_{1}(\lambda\sigma^{-1}m_{1}([\alpha]))\end{split}$
@@
In the set-up of § 2.1 let $y_{0}\in Y$, let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$ and let $(f:X\to Y,x_{0})$ be a pointed cover of $(Y,y_{0})$ of degree $d$ branched in $D$. The following conditions are equivalent:
  1. (i)
@@
    (3) $\begin{split}\varepsilon(\lambda m([\alpha]))&=\varepsilon(\lambda)\alpha\quad\text{for}\quad\forall\lambda\in\Lambda,\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})\\ \varepsilon(\lambda_{0})&=x_{0}\end{split}$
  2. (ii)
@@
The conditions (i) and (ii) are equivalent since the map $\pi_{1}(Y\setminus D,y_{0})\to f^{-1}(y_{0})$ defined by $[\alpha]\mapsto x_{0}\alpha$ induces a $\pi_{1}(Y\setminus D,y_{0})$-equivariant bijection between the set of right cosets \${}_{\Gamma_{x_{0}}}\backslash\pi_{1}(Y\setminus D,y_{0})$ and $f^{-1}(y_{0})$, $m$ induces a $\pi_{1}(Y\setminus D,y_{0})$-equivariant bijection between \${}_{\Gamma_{x_{0}}}\backslash\pi_{1}(Y\setminus D,y_{0})$ and \${}_{G(\lambda_{0})}\backslash G\cong\Lambda$. Under these bijections $\lambda_{0}\mapsto G(\lambda_{0})\mapsto\Gamma_{\lambda_{0}}\mapsto x_{0}$.
@@
In the set-up of § 2.1 let $N(\lambda_{0})=\{\sigma\in N_{S(\Lambda)}(G)|\lambda_{0}\sigma=\lambda_{0}\}$. Consider the left action of $N(\lambda_{0})$ on $H^{G}_{n}(Y,y_{0})$ defined by $\sigma\ast(D,m)=(D,\sigma m\sigma^{-1})$. We denote by $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ the quotient set
@@
Let $G(\lambda_{0})=H$. Let $\rho:C\to C/H$ be the quotient morphism. One has $\Lambda\times^{G}C\cong C/H$. In fact, let us choose for every $\lambda\in\Lambda$ an element $a_{\lambda}\in G$ such that $\lambda=\lambda_{0}a_{\lambda}$, let $a_{\lambda_{0}}=1$. The morphisms $C\to\Lambda\times^{G}C$ and $\Lambda\times C\to C/H$ defined respectively by
(5) $z\mapsto\pi(\lambda_{0},z),\quad(\lambda,z)\mapsto\rho(a_{\lambda}z)$
@@
Let $y\in Y\setminus D$ and let $z\in p^{-1}(y)$. The map $\Lambda\to f^{-1}(y)$ defined by $\lambda\mapsto\pi(\lambda,z)$ is bijective, so $f$ is unbranched at every $y\in Y\setminus D$. The monodromy homomorphism $m=m_{z_{0}}$ satisfies Condition (1), so every $[\gamma_{i}]$ acts on the fiber $f^{-1}(y_{0})$ as a nontrivial permutation, $i=1,\ldots,n$. Therefore the branch locus of $f:X\to Y$ equals $D$. We see that $(f:X\to Y,x_{0})$ satisfies the conditions of Proposition 2.8(i) with $(\varepsilon,m)$, so its monodromy invariant is $(D,m^{N(\lambda_{0})})$. ∎
@@
The map $[f:X\to Y,x_{0}]\mapsto(D,m^{N(\lambda_{0})})$ establishes a bijective correspondence between the set of equivalence classes of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points and the set $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$.
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y$. Let $f:X\to Y$ be a cover branched in $D$. The following conditions are equivalent.
  1. (i)
@@
    $\varepsilon(\lambda m([\alpha]))=\varepsilon(\lambda)\alpha\quad\text{for}\quad\forall\lambda\in\Lambda,\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$
  2. (ii)
@@
$m_{2}([\alpha])=\sigma m_{1}([\gamma\cdot\alpha\cdot\gamma^{-}])\sigma^{-1}\quad\text{for}\quad\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{2})$
@@
$\varepsilon_{1}(\lambda m_{1}([\beta]))=\varepsilon(\lambda m([\gamma\cdot\beta\cdot\gamma^{-}]))\gamma=\varepsilon(\lambda)\gamma\cdot\beta\cdot\gamma^{-}\cdot\gamma=\varepsilon_{1}(\lambda)\beta$
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y$. Given two points $y_{1},y_{2}\in Y\setminus D$ and two homomorphisms $m_{1}:\pi_{1}(Y\setminus D,y_{1})\to G$ and $m_{2}:\pi_{1}(Y\setminus D,y_{2})\to G$, $G\subset S(\Lambda)$, we write $m_{1}\sim_{N}m_{2}$ if there is a path $\gamma:I\to Y\setminus D$ with $\gamma(0)=y_{1}$, $\gamma(1)=y_{2}$ and an element $\sigma\in N_{S(\Lambda)}(G)$, such that $m_{2}([\alpha])=\sigma m_{1}([\gamma\cdot\alpha\cdot\gamma^{-}])\sigma^{-1}$ for $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{2})$. This is a relation of equivalence. Given a homomorphism $m:\pi_{1}(Y\setminus D,y)\to G$ we denote by $[m]$ its equivalence class. It is clear that if $m$ is an epimorphism which satisfies Condition (1) relative to the base point $y$ every other homomorphism of $[m]$ has these properties relative to its base point. Suppose a cover $f:X\to Y$ branched in $D$ satisfies Condition (i) of Proposition 2.15 with $(\varepsilon,m)$ for some $y_{0}\in Y$. Then the set of all possible epimorphisms $m_{1}:\pi_{1}(Y\setminus D,y)\twoheadrightarrow G$, $y\in Y\setminus D$ as in Proposition 2.15(ii) equals the equivalence class $[m]$ (apply Proposition 2.15 to $f_{1}=f=f_{2}$, $y_{1}=y_{0}$, $y_{2}=y$). We denote by $[m]_{f}$ the equivalence class with respect to $\sim_{N}$ determined by $f:X\to Y$.

### v4

@@ Abstract.
This version of the article has been accepted for publication, after peer review, in Atti Accad. Naz. Lincei Rend. Lincei Mat. Appl., https://doi.org/10.4171/RLM/1061,
©  2025 Accademia Nazionale dei Lincei.
@@ 1. Introduction
The covers $f:X\to Y$ with restricted monodromy group and the related Galois covers $p:C\to Y$ yield polarized abelian varieties isogenous to abelian subvarieties of the Jacobian variety $J(X)$ [6, 17, 26, 4]. Given a smooth, proper family of such covers one obtains a morphism of its parameter variety to a certain moduli space of polarized abelian varieties, which may be studied by means of the associated variation of the Hodge structures of weight one [13, 14, 3]. This approach to the moduli of polarized abelian varieties was used in [18, 19] where the unirationality of $\mathcal{A}_{3}(1,1,d)$ and $\mathcal{A}_{3}(1,d,d)$ with $2\leq d\leq 4$ was proved by means of families of simply ramified covers of elliptic curves branched in 6 points. The Prym–Tyurin morphism from a Hurwitz space to $\mathcal{A}_{6}$ was used in [1] to prove that every sufficiently general principally polarized abelian variety of dimension 6 is isomorphic to a Prym–Tyurin variety of a cover of $\mathbb{P}^{1}$ of degree 27, branched in 24 points, with monodromy group the Weyl group $W(E_{6})\subset S_{27}$.
@@
We assume the base field is $\mathbb{C}$. Algebraic varieties are reduced, separated, possibly reducible schemes of finite type, points are closed points. Fiber products and pullbacks are those defined in the category of schemes over $\mathbb{C}$. A cover $f:X\to Z$ of algebraic varieties is a finite, surjective morphism. If $Y$ is a smooth, projective, irreducible curve we denote by $Y^{(n)}$ the $n$-th symmetric power of $Y$ and by $Y^{(n)}_{\ast}$ the subset of $Y^{(n)}$ consisting of effective divisors of $Y$ of degree $n$ without multiple points. We identify such divisors with their support. If $\Omega\subset Y^{(n)}$, then $\Omega_{\ast}=\Omega\cap Y^{(n)}_{\ast}$. Given an algebraic variety $(X,\mathcal{O}_{X})$ the canonically associated complex space is denoted by $(X^{an},\mathcal{O}_{X^{an}})$ [28]. Its topological space is denoted by $|X^{an}|$. We assume that the homotopy of paths leaves the end points fixed [24, Chapter 2, §2]. The set of paths homotopic to $\alpha$ is denoted by $\alpha]$. The product of the paths $\alpha$ and $\beta$ is denoted by $\alpha\cdot\beta$ and it equals the path $\gamma$, where $\gamma(t)=\alpha(2t)$ if $t\in[0,\frac{1}{2}]$, $\gamma(t)=\beta(2t-1)$ if $t\in[\frac{1}{2},1]$. Given a covering space $p:M\to N$ of the topological space $N$, the map $p$ is called topological covering map. Lifting a path $\alpha$ of $N$ from an initial point $z\in M$ the terminal point is denoted by $z\alpha$.
@@ 2. Parameterization of covers with a fixed monodromy group
Throughout the paper $Y$ is a smooth, projective, irreducible curve of genus $g\geq 0$, $n$ is a positive integer, $G$ is a finite group, which can be generated by 2g+n12g+n-1 elements, $\Lambda$ is a finite set, $|\Lambda|=d$, on which $G$ acts on the right faithfully and transitively, we choose an element $\lambda_{0}\in\Lambda$ as a marked element, $G(\lambda_{0})=\{g\in G|\lambda_{0}g=\lambda_{0}\}$. We identify $G$ with its image in the symmetric group $S(\Lambda)\cong S_{d}$, where $S(\Lambda)$ acts on the right.
@@
A necessary condition for two covers $f:X\to Y$ and $f_{1}:X_{1}\to Y$ to be equivalent is that they have the same branch locus $D$. Let $y_{0}\in Y\setminus D$. Let $Y^{an}$ be the Riemann surface associated with $Y$. Then $\pi_{1}(Y^{an}\setminus D,y_{0})$ acts on the right on $f^{-1}(y_{0})$ by $x\cdot[\alpha]=x\alpha$ for $\forall x\in f^{-1}(y_{0})$ and $\forall[\alpha]\in\pi_{1}(Y^{an}\setminus D,y_{0})$, and similarly it acts on the right on $f_{1}^{-1}(y_{0})$. Then $f$ is equivalent to $f_{1}$ if and only if there exists a $\pi_{1}(Y^{an}\setminus D,y_{0})$-equivariant bijection $\mu:f^{-1}(y_{0})\to f_{1}^{-1}(y_{0})$. In fact let $X^{\prime}=f^{-1}(Y\setminus D)$, $X_{1}^{\prime}=f_{1}^{-1}(Y\setminus D)$. Then such a bijection yields a covering biholomorphic map $X^{\prime an}\to X^{\prime an}_{1}$ over $Y\setminus D$, which extends to a covering biholomorphic map $X^{an}\to X^{an}_{1}$ over $Y^{an}$. Furthermore this map equals $h^{an}$, where $h:X\to X_{1}$ is a covering isomorphism over $Y$. Abusing notation we will write $\pi_{1}(Y\setminus D,y_{0})$ instead of $\pi_{1}(Y^{an}\setminus D,y_{0})$.
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$. Let $\overline{U}_{1},\ldots,\overline{U}_{n}$ be embedded closed disks in $Y\setminus y_{0}$ which are disjoint and such that $b_{i}\in U_{i}$ for $\forall i$, where $U_{i}$ is the interior of $\overline{U}_{i}$. For every $i=1,\ldots,n$ let us choose a path $\eta_{i}:I\to Y\setminus\cup_{j=1}^{n}U_{j}$ such that $\eta_{i}(0)=y_{0}$, $\eta_{i}(1)\in\partial\overline{U}_{i}$ and let $\gamma_{i}:I\to Y\setminus D$ be the closed path which starts at $y_{0}$, travels along $\eta_{i}$, then makes a counterclockwise loop along $\partial\overline{U}_{i}$ and returns back to $y_{0}$ along $\eta_{i}^{-}$. Let $f:X\to Y$ be a $d$-sheeted cover of $Y$ unbranched at $y_{0}$. Let $X^{\prime}=X\setminus f^{-1}(D)$, $f^{\prime}=f|_{X^{\prime}}$. The branch locus of $f$ equals $D$ if and only if $f^{\prime}:X^{\prime}\to Y\setminus D$ is unramified and the monodromy $m:\pi_{1}(Y\setminus D,y_{0})\to S(f^{-1}(y_{0}))$ defined by $x\cdot m([\alpha])=x\alpha$ for $\forall x\in f^{-1}(y_{0})$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$ satisfies the condition
@@
Let $m:\pi_{1}(Y\setminus D,y_{0})\to S_{d}$ be a homomorphism and let $G=\operatorname{Im}(m)$. Suppose that $m$ satisfies Condition (1). If one chooses in a different way $\overline{U}_{1},\ldots,\overline{U}_{n}$ and $\eta_{1},\ldots,\eta_{n}$ so as to satisfy the above conditions, then for every $i\in[1,n]$ the new element $\gamma^{\prime}_{i}]$ belongs to the conjugacy class of $\gamma_{i}]$ in $\pi_{1}(Y\setminus D,y_{0})$, so $m([\gamma^{\prime}_{i}])$ belongs to the conjugacy class of $m([\gamma_{i}])$ in $G$. Therefore Condition (1) for $m$ is independent of the choice of $\overline{U}_{1},\ldots,\overline{U}_{n}$ and that of $\eta_{1},\ldots,\eta_{n}$.
@@
Let $G\subset S(\Lambda)$ be as in §2.1. Let $N_{S(\Lambda)}(G)=\{\sigma\in S(\Lambda)|\sigma G\sigma^{-1}=G\}$. Let $f:X\to Y$ and $f_{1}:X_{1}\to Y$ be two covers with the same branch locus $D$. Let $y_{0}\in Y\setminus D$. Suppose that there are bijections $\varepsilon:\Lambda\to f^{-1}(y_{0})$, $\varepsilon_{1}:\Lambda\to f_{1}^{-1}(y_{0})$ and epimorphisms $m:\pi_{1}(Y\setminus D,y_{0})\to G$, $m_{1}:\pi_{1}(Y\setminus D,y_{0})\to G$ such that $\varepsilon(\lambda m([\alpha]))=\varepsilon(\lambda)\alpha$ and $\varepsilon_{1}(\lambda m_{1}([\alpha]))=\varepsilon_{1}(\lambda)\alpha$ for $\forall\lambda\in\Lambda$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$. Then $f$ is equivalent to $f_{1}$ if and only if there exists a $\sigma\in N_{S(\Lambda)}(G)$ such that $m_{1}=\sigma m\sigma^{-1}$. The covering isomorphism $h:X\to X_{1}$ is unique if and only if the centralizer of $G$ in $S(\Lambda)$ is trivial, $Z_{S(\Lambda)}(G)=\{1\}$.
@@
The covers $f:X\to Y$ and $f_{1}:X_{1}\to Y$ are equivalent if and only if there exists a $\pi_{1}(Y\setminus D,y_{0})$-equivariant bijection $\mu:f^{-1}(y_{0})\to f_{1}^{-1}(y_{0})$ (cf. §2.3). Let $\mu$ be such a bijection. Let $s:\Lambda\to\Lambda$ be the bijection, which makes the following diagram commutative:
[…]
@@
$\begin{split}&\mu(x\alpha)=\mu(\varepsilon(\lambda)\alpha)=\mu(\varepsilon(\lambda m([\alpha])))=\varepsilon_{1}(\lambda m([\alpha])\sigma^{-1})\\ &\mu(x)\alpha=\mu(\varepsilon(\lambda))\alpha=\varepsilon_{1}(\lambda\sigma^{-1})\alpha=\varepsilon_{1}(\lambda\sigma^{-1}m_{1}([\alpha])).\end{split}$
@@
In the setup of §2.1 let $y_{0}\in Y$, let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$ and let $(f:X\to Y,x_{0})$ be a pointed cover of $(Y,y_{0})$ of degree $d$ branched in $D$. The following conditions are equivalent:
  1. (1)
@@
    (3) $\begin{split}\varepsilon(\lambda m([\alpha]))&=\varepsilon(\lambda)\alpha\quad\text{for}\quad\forall\lambda\in\Lambda,\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})\\ \varepsilon(\lambda_{0})&=x_{0}.\end{split}$
  2. (2)
@@
The conditions (i) and (ii) are equivalent since the map $\pi_{1}(Y\setminus D,y_{0})\to f^{-1}(y_{0})$ defined by $\alpha]\mapsto x_{0}\alpha$ induces a $\pi_{1}(Y\setminus D,y_{0})$-equivariant bijection between the set of right cosets $_{\Gamma_{x_{0}}}\backslash\pi_{1}(Y\setminus D,y_{0})$ and $f^{-1}(y_{0})$, $m$ induces a $\pi_{1}(Y\setminus D,y_{0})$-equivariant bijection between $_{\Gamma_{x_{0}}}\backslash\pi_{1}(Y\setminus D,y_{0})$ and ${G(\lambda_{0})}\backslash G\cong\Lambda$. Under these bijections $\lambda_{0}\mapsto G(\lambda_{0})\mapsto\Gamma_{\lambda_{0}}\mapsto x_{0}$.
@@
In the setup of §2.1 let $N(\lambda_{0})=\{\sigma\in N_{S(\Lambda)}(G)|\lambda_{0}\sigma=\lambda_{0}\}$. Consider the left action of $N(\lambda_{0})$ on $H^{G}_{n}(Y,y_{0})$ defined by $\sigma\ast(D,m)=(D,\sigma m\sigma^{-1})$. We denote by $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ the quotient set
@@
Let $G(\lambda_{0})=H$. Let $\rho:C\to C/H$ be the quotient morphism. One has $\Lambda\times^{G}C\cong C/H$. In fact, let us choose for every $\lambda\in\Lambda$ an element $a_{\lambda}\in G$ such that $\lambda=\lambda_{0}a_{\lambda}$, let $a_{\lambda_{0}}=1$. The morphisms $C\to\Lambda\times^{G}C$ and $\Lambda\times C\to C/H$, defined by
(5) $z\mapsto\pi(\lambda_{0},z)\quad\text{and}\quad(\lambda,z)\mapsto\rho(a_{\lambda}z),$
@@
Let $y\in Y\setminus D$ and let $z\in p^{-1}(y)$. The map $\Lambda\to f^{-1}(y)$ defined by $\lambda\mapsto\pi(\lambda,z)$ is bijective, so $f$ is unbranched at every $y\in Y\setminus D$. The monodromy homomorphism $m=m_{z_{0}}$ satisfies Condition (1), so every $\gamma_{i}]$ acts on the fiber $f^{-1}(y_{0})$ as a nontrivial permutation, $i=1,\ldots,n$. Therefore the branch locus of $f:X\to Y$ equals $D$. We see that $(f:X\to Y,x_{0})$ satisfies the conditions of Proposition 2.8(i) with $(\varepsilon,m)$, so its monodromy invariant is $(D,m^{N(\lambda_{0})})$. ∎
@@
The map $f:X\to Y,x_{0}]\mapsto(D,m^{N(\lambda_{0})})$ establishes a bijective correspondence between the set of equivalence classes of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points and the set $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$.
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y$. Let $f:X\to Y$ be a cover branched in $D$. The following conditions are equivalent:
  1. (1)
@@
    $\varepsilon(\lambda m([\alpha]))=\varepsilon(\lambda)\alpha\quad\text{for}\quad\forall\lambda\in\Lambda,\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0}).$
  2. (2)
@@
$m_{2}([\alpha])=\sigma m_{1}([\gamma\cdot\alpha\cdot\gamma^{-}])\sigma^{-1}\quad\text{for}\quad\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{2}).$
@@
$\varepsilon_{1}(\lambda m_{1}([\beta]))=\varepsilon(\lambda m([\gamma\cdot\beta\cdot\gamma^{-}]))\gamma=\varepsilon(\lambda)\gamma\cdot\beta\cdot\gamma^{-}\cdot\gamma=\varepsilon_{1}(\lambda)\beta.$
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y$. Given two points $y_{1},y_{2}\in Y\setminus D$ and two homomorphisms $m_{1}:\pi_{1}(Y\setminus D,y_{1})\to G$ and $m_{2}:\pi_{1}(Y\setminus D,y_{2})\to G$, $G\subset S(\Lambda)$, we write $m_{1}\sim_{N}m_{2}$ if there is a path $\gamma:I\to Y\setminus D$ with $\gamma(0)=y_{1}$, $\gamma(1)=y_{2}$ and an element $\sigma\in N_{S(\Lambda)}(G)$, such that $m_{2}([\alpha])=\sigma m_{1}([\gamma\cdot\alpha\cdot\gamma^{-}])\sigma^{-1}$ for $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{2})$. This is a relation of equivalence. Given a homomorphism $m:\pi_{1}(Y\setminus D,y)\to G$ we denote by $[m]$ its equivalence class. It is clear that if $m$ is an epimorphism which satisfies Condition (1) relative to the base point $y$ every other homomorphism of $[m]$ has these properties relative to its base point. Suppose a cover $f:X\to Y$ branched in $D$ satisfies Condition (i) of Proposition 2.15 with $(\varepsilon,m)$ for some $y_{0}\in Y$. Then the set of all possible epimorphisms $m_{1}:\pi_{1}(Y\setminus D,y)\twoheadrightarrow G$, $y\in Y\setminus D$ as in Proposition 2.15(ii) equals the equivalence class $[m]$ (apply Proposition 2.15 to $f_{1}=f=f_{2}$, $y_{1}=y_{0}$, $y_{2}=y$). We denote by $m]_{f}$ the equivalence class with respect to $\sim_{N}$ determined by $f:X\to Y$.